DETERMINATION OF THE STRUCTURE OF METHYLENE CYCLOPENTANE AND THE ARGON METHYLENE CYCLOPENTANE VAN DER WAALS COMPLEX

\underline{ANDREA J. MINEI}, WALLACE C. PRINGLE, and STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

Rotational spectra of methylene cyclopentane (MCP) and the argon methylene cyclopentane van der Waals complex were studied using pulsed-jet Fabry-Perot Fourier transform microwave spectroscopy. Spectra of the normal isotopologue along with those of the natural abundance ^{13}C isotopologues for both the monomer and the complex were observed and assigned in the frequency region of 5-20 GHz. The substitution structure of the monomer has the two methylene carbons on the C_{2} axis and the four other ring carbons confirm the twisted conformation. The two ring carbons attached to the ring methylene carbon are 0.14 Å above and 0.14 Å below the methylene plane. The two carbons at the back of the ring are 0.25 Å below and 0.25 Å above the plane of the methylene group. For the argon MCP complex, a- and b-type transitions were observed and the rotational constants were determined to be A = 2472.4022(4), B = 1088.2403(2),and C = 937.9203(2) MHz. The argon atom is situated 3.6 Å above the methylene group plane and 0.1 Å behind the methylene ring carbon. Additionally, the argon atom is 0.8 Å across the ring from the center of mass of MCP and is above the ring carbon that is 0.14 Å below the plane. An investigation of large amplitude cross ring motion of the argon atom as well as comparison to previous ring complexes studied in this laboratory will be presented.